if a boy is 4years older than his sister and there combined age is the sum of 14 then how old is the boy? what is the girl age?
Answer by checkley77(12844)

X+X+4=14
2X=14-4
2X=10
X=10/2
X=5 FOR THE AGE OF THE SISTER.
5+4=9 FOR THE AGE OF THE BOY.
PROOF:
5+9=14
14=14